OpenStudy (jb1625):

What property is shown in the equation? 3ac = 3ca A. Identity Property of Multiplication B. Reciprocal Property of Multiplication C. Zero Property of Multiplication D. Commutative Property of Multiplication

OpenStudy (anonymous):

Identity Property of Multiplication: a*1=a Reciprocal Property of Multiplication: 1/a*a=1 Zero Property of Multiplication: 0*a=0 Commutative Property of Multiplication: ab=ba

OpenStudy (jb1625):

So what is the answer?

OpenStudy (anonymous):

D

OpenStudy (jb1625):

Ok Thank You!! :)

OpenStudy (anonymous):

np :)